OpenStudy (anonymous):

Two numbers are in the ratio of 2 to 3. If the smaller number is 18, the larger number is _____.

OpenStudy (anonymous):

X * 2/3 = 18 solve for X X = 18 * 3/2 X = 27

OpenStudy (anonymous):

18:27 ---> 2:3

OpenStudy (anonymous):